OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (3, 0) and a directrix at x = -3. @aum

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@ganeshie8 @mathmate @mathstudent55

OpenStudy (aum):

This is a horizontal parabola that opens to the right. Any point on the parabola is equidistant from the focus and the directrix: (x-3)^2 + (y-0)^2 = (x - (-3))^2 (x-3)^2 + y^2 = (x+3)^2 y^2 = (x+3)^2 - (x-3)^2 y^2 = (x+3+x-3)(x+3-x+3) y^2 = 2x * 6 = 12x x = 1/12 * y^2

OpenStudy (anonymous):

thank you

OpenStudy (aum):

yw.